Find the point of intersection of the graphs of x-y=-18 and x+y=-9.: Find the point of intersection of the graphs of x-y=-18 and x+y=-9.
Answer by mangopeeler07(448)

All you have to do to solve this problem is a little elimination. Since in one equation 1y is positive and in the other 1y is negative, it is easy to eliminate. Simply line x-y=-18 up to x+y=-9 and add the expression and their answers. You should have gotten 2x=-27. Then divide each side by two and get x=-13.5. Then plug x in the first equation and end up with y=4.5. -13.5-y=-18. Your solutions for x and y are the coordinates of the point of intersection of the two lines.